the population P of a certain culture is expected to be given by a mdoel P=100e^(rt) where r is a constant to be determined and t is a number of days since the original population of 100 was established. Find the r value if the population is expected to reach 200 in 3 days.

Answer by Nate(3500)

P = 100e^(rt)
ln(P/100) = rt
ln(P/100)/t = r
ln(2)/3 = r = 0.23105 = 23%